When 50 ml water sample titrated against 0.02 N HCl acid using phenolphthalein



indicator, the titration reading was 24.2 ml. After boiling 100 ml of the same sample using



methyl orange indicator, the titration value was 28.2 ml. Calculate permanent and total hardness.

1
Expert's answer
2022-01-16T22:56:02-0500



Hardness "(mg" of "CaCO_3/L)=\frac{A.N.50000}{ML\>sample}"


Where "A=" ml of standard acid used to titrate the sample

"N=" Normality of the standard acid


Temporary hardness "=\frac{24.2×0.02×50000}{50}"


"=484\>ppm"


Permanent Hardness"=\frac{28.2×0.02×50000}{100}"


"=282\>ppm"


Total Hardness "=766\>ppm"